NON-CYCLIC ISOMERS OF (H_{2}O)_{4} IN HELIUM NANODROPLETS: INFRARED SPECTROSCOPY AND AB INITIO CALCULATIONS

\underline{S. D. FLYNN}, A. M. MORRISON, T. LIANG, and G. E. DOUBERLY, DEPARTMENT OF CHEM-ISTRY, UNIVERSITY OF GEORGIA, ATHENS, GEORGIA 30602-2556; S. S. XANTHEAS, CHEMICAL AND MATERIALS SCIENCES DIVISION, PACIFIC NORTHWEST NATIONAL LABORATORY, 906 BATTELLE BOULEVARD, MS K1-83, RICHLAND, WASHINGTON 99352.

Water clusters are assembled via the sequential pick-up of water molecules by helium nanodroplets. Unlike previous infrared spectroscopy experiments of water clusters in helium droplets,^{a} one or two Neon atoms are added to the droplets prior to water pick-up. The upstream pick-up of a Neon atom results in several new bands in the infrared spectrum in addition to the bands that correspond to the water monomer, dimer and larger cyclic water complexes. The new spectral features are determined to be signatures of a (H_{2}O)_{4} cluster on the basis of the pick-up cell pressure dependence of the band intensities. A dc electric field is applied to the laser droplet beam interaction region, and these clusters are determined to be polar with permanent dipole moments between 2 and 3 Debye. On the basis of comparisons to CCSD(T) anharmonic frequency calculations, the new bands are assigned to OH stretch vibrations of a non-cyclic 3+1 cluster, which corresponds to a water molecule hydrogen bonded to a trimer ring. The presence of the Neon atom substantially affects the barrier to ring insertion of the fourth water molecule into a preformed cyclic trimer complex. In contrast, no new bands corresponding to open (non-cyclic) trimers or pentamers are observed.

aK. Nauta and R. E. Miller, Science 287, 293 (2000).